December 3, 2015 Drift Chamber meeting
- Instructions for Bluejeans meeting connection
- Meeting ID: 290664653
- To join via a Web Browser, go to the page  https://bluejeans.com/290664653.
- CDC update (Mike, Simon)
- FDC update
- FDC check with cosmics:  (Luke)
- Gain drop with 55Fe: ,  (Lubomir)
- Instructions for the TRD test in the fall 2015 run, radiator in front of TRD
- Electronics (Fernando, Chris, Nick)
- Engineering (Dave)
Participants: Naomi, Mike (CMU), Sean (NU), Cody, Luke, Dave, Nick, Simon, Beni, Eugene (at the end), and Lubomir (JLab).
- Mike showed plots with time-to-distance deformation corrections now WITH magnetic field. They look close to the Garfield predictions. Next step is to apply the corrections and look at the beam data also for different runs.
- Cody fixed several bugs in the firmware and updated the fADCs in the CDC crates only. The issues were found by Naomi: end of window algorithm fails if no peak was found, issues with the overflow bit when summing; and other. The plan is to have a cosmic run today/tonight with CDC to check the fixes. Cody also swapped two CDC modules. Cody is out starting this Monday.
- Luke looked at the cosmic data with the FDC and updated the spreadsheet (linked above) with the problematic channels: green - channels that were not working before and now are OK, red - new bad channels, all the rest no change in the status. The green is more than the red meaning we have lower percentage of bad channels, which is of the order of 0.5%.
- Looking at the results with 55Fe source on the spare package Lubomir found that the gain at the wire closest to the source is factor of 2 lower than the gain on the side wires (first plot with 1700V/-0V, second with 1600V/-400V Ar/CO2 90/10). The averaged charge for each wire is normalized by moving the source away from the front window. The gain drops due to the high rates. Using a scope and lowering the threshold till the rate stopped depending on it, we measured with Luke rates of ~40kHz. In case of MIP in the standard cells we expect the same gain drop by a factor of two to happen at ~100kHz (ratio of primary electrons for (6keV photons)/(1cm MIP) is 230/90).
- The TRD test is ready to go (just need beam) - follow the link above. Sergey and Lubomir installed radiators with different thicknesses in front of the spare package (see picture above).
- Nick is going to look at the last CDC LV module. It had 4 channels (out of 8) connected and was OK, but then we connected 3 extra channels for the spare FDC package and showed some issues. The plan is to replace the module.
- Lubomir will be out of the country next two weeks, accessible via e-mails. Next DC meeting will be in January, unless there is an additional announcement.